OpenStudy (anonymous):

how would i integrate 5xe^(-x)

OpenStudy (anonymous):

Integration by parts

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

\[ \int f\cdot g' = fg-\int f'\cdot g \]

OpenStudy (anonymous):

In this example it's easiest to set \(g'=e^{-x}, f=x\), you can just pull the 5 outside of the integral.

OpenStudy (anonymous):

what is the correct answer?

OpenStudy (lgbasallote):

in choosing f you follow this order of priority 1. logarithms 2. inverse trig functions 3. algebraic expressions 4. trig functions 5. exponential functions

OpenStudy (anonymous):

What are you getting as an answer? I'll let you know whether it's correct or not.

OpenStudy (anonymous):

i got -xe^(-x)+e^(-x)+c

OpenStudy (anonymous):

Don't forget about the 5 that you originally pulled out.

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Also, to verify an answer, just take the derivative and see if you get what you started with :)

OpenStudy (anonymous):

ok i will do that right now

OpenStudy (anonymous):

Put the 5 back in on each term, though, and yes that is correct. Still worth deriving to double-check though.

OpenStudy (anonymous):

yay thanks. so that means my answer should have been -5xe^(-x)+5e^(-x) right?

OpenStudy (anonymous):

Actually I think it's -5xe^(-x)-5e^(-x). Triple negative on the second term, one from the integration by parts formula, one from g, and one from integrating g. Sorry, doing this all in my head not with pen and paper.

OpenStudy (anonymous):

thank you